The average pre-school cost for tuition fees last year was ₱17,250. The following tear, 20 institutions had a mean of ₱16,150 and a standard deviation of ₱2,250. Is there sufficient evidence at a = 0.10 level of significance to conclude that the mean cost has increased?


1
Expert's answer
2022-03-16T16:33:15-0400

Null hypothesis H0: "\mu=17250"

Alternative hypothesis Ha: "\mu>17250"


Sample size "n=20"

Sample mean "\bar X=16150"

Sample standard deviation "\sigma =2250"

Significance level "\alpha=0.10"


Since the alternative hypothesis is that "\mu" is greater than this is a right tailed hypothesis test.


Since the sample size is less than 30, we use the T test


"df=n-1=20-1=19"


Since "\alpha=0.10" "T=1.328"


"t_c=\dfrac{\bar X-\mu_0}{S/\sqrt n}" "=\dfrac{16150-17250}{2250/\sqrt {20}}"


"t_c=-2.186"


Since "t_c<T" we reject the hypothesis.


Since the null hypothesis is rejected, there is sufficient evidence to conclude that the population mean cost has increased at 0.10 level of significance.